# Semi-Blind Strategies for MMSE Channel Estimation Utilizing Generative Priors

### v1

@@ IV-A GMM-based Semi-blind Channel Estimation
where $p(k)$, $\bm{\mu}_{k}$ and ${\bm{C}}_{k}$ are the mixing coefficients, means and covariance matrices of the $k$-th GMM component, respectively. As we are considering wireless channels, the mean of each component is set to $\bm{\mu}_{k}=\bm{0}$, cf. [39]. The fitting of the components in (40) is accomplished with the well-known EM algorithm [41] based on a set =$\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ of $T$ channel samples as training data. Based on the formulation in (40) the conditional PDF given $k$ is
@@
Thus, in the case of a GMM we have a discrete latent variable, which helps us parameterize the CME. The resulting semi-blind subspace GMM can be formulated as
@@
Algorithm 1 Subspace GMM Channel Estimator
Offline Training Phase
1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the GMM with the EM algorithm, cf. [30]
3:
4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
8:for $j=1,\dots,J$ do
9:     ${\bm{y}}^{\prime}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:     for $k=1,\dots,K$ do
12:     end for
14:end for
15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
Algorithm 2 Projected GMM Channel Estimator
Offline Training Phase
1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the GMM with the EM algorithm, cf. [30]
3:
4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
8:for $j=1,\dots,J$ do
9:     $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:     for $k=1,\dots,K$ do
12:     end for
14:end for
15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
@@
IV-B VAE-based Semi-blind Channel Estimation
To learn the unknown distribution $f_{\bm{h}}({\bm{h}})$ using a VAE, we lower bound the parameterized likelihood $p_{\bm{\theta}}({\bm{h}})$ using the evidence-lower bound (ELBO). To formulate the ELBO, the variational distributions $q_{\bm{\phi}}({\bm{z}}\mid{\bm{y}}^{\prime})$ and $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$ are introduced, which approximate $p({\bm{z}}\mid{\bm{y}}^{\prime})$ and $p({\bm{z}}\mid\tilde{{\bm{y}}})$, respectively. In contrast to the GMM, the used subspace $\mathrm{range}({\bm{V}})$ is unknown to the encoder of the VAE making $p({\bm{z}}\mid{\bm{y}}^{\prime})$ difficult to learn. Additionally, the dimension of the encoder input would depend on the number of users in the system. Thus, we propose to approximate both posteriors $p({\bm{z}}\mid{\bm{y}}^{\prime})$ and $p({\bm{z}}\mid\tilde{{\bm{y}}})$ with $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$. A version of the ELBO for this case, which is accessible, can be written as [42]
@@ IV-B VAE-based Semi-blind Channel Estimation
where $\mathbb{E}_{q_{\bm{\phi}}}[\cdot]=\mathbb{E}_{q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})}[\cdot]$ is the expectation over the variational distribution $q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$. The second term in (48) is the Kullback-Leibler (KL) divergence
@@
$q_{\bm{\phi}}({\bm{z}}\mid\tilde{{\bm{y}}})$ $=\mathcal{N}(\bm{\mu}_{\bm{\phi}}(\tilde{{\bm{y}}}),\mathrm{diag}(\bm{\sigma}^{2}_{\bm{\phi}}(\tilde{{\bm{y}}}))).$
The resulting semi-blind VAE structure is shown in Fig. 2. In the case of a ULA or URA at the BS the channel covariance matrix is either Toeplitz or block-Toeplitz, respectively. As shown in [39], the conditional covariance matrix at the output of the VAE preserves this structure. Thus, we parameterize the output covariance matrix as
@@
where ${\bm{Q}}={\bm{Q}}_{M}$ or ${\bm{Q}}={\bm{Q}}^{\prime}_{N_{v}}\otimes{\bm{Q}}^{\prime}_{N_{h}}$, respectively, where ${\bm{Q}}_{M}$ is a DFT matrix of size $M$ resulting in a circulant approximation, cf. [32], and ${\bm{Q}}^{\prime}_{N_{x}}$ contains the first $N_{x}$ columns of the $2N_{x}\times 2N_{x}$ DFT matrix resulting in a block-Toeplitz parameterization, cf. [43]. Further, for the (block-)Toeplitz parameterization we can set $\bm{\mu}_{\bm{\theta}}({\bm{z}})=\bm{0}$, cf. [39].
After successfully training the VAE, the output is a local parameterization of $f_{\bm{h}}({\bm{h}})$ as conditionally Gaussian
${\bm{h}}\mid{\bm{z}}\sim p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}}).$ (52)
As analyzed in [32] it is a reasonable approximation to set
Based on this parameterization we can formulate the semi-blind VAE-based estimators as
[…]
and
[…]
×$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}).$ (55)
The respective estimators are summarized in Algorithm 3 and Algorithm 4. For a more detailed introduction into the VAE framework and its usage for parameterization of the CME we refer the reader to [32].
@@
1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
@@
4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
@@
12:           ×$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}})$
@@
1:Training dataset =$\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
@@
4:${\bm{Y}}=[{\bm{y}}(1),\dots,{\bm{y}}(N)]$, ${\bm{P}}$, $\sigma^{2}$
5:$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}\leftarrow\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}$
6:$\hat{{\bm{V}}}\leftarrow J$ dominant eigenvectors of $\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}$
7:${\bm{Y}}_{p}=[{\bm{y}}_{p,1},\dots,{\bm{y}}_{p,J}]\leftarrow{\bm{Y}}^{\prime}_{p}{\bm{P}}^{\mathrm{H}}$
@@
12:           ×$\times(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
@@
IV-C Maximum Likelihood Subspace Estimation
After introducing the methods utilizing the additional subspace information provided by $\mathrm{range}({\bm{V}})$ to enhance the CSI estimation quality, let us consider the estimation of such a subspace. As the received data symbols are also transmitted over the same channel, they can be used to estimate the subspace containing all user channels.
To this end, let us reconsider the ML estimate of ${\bm{H}}$ in (13). Instead of directly optimizing on this ML formulation as done in [16], which generally does not result in the MMSE, we only take this log-likelihood formulation as an intermediate step to estimate the subspace $\mathrm{range}({\bm{V}})$. First, let us consider the right term of the objective function in (13). We can then reformulate the problem by again solving for ${\bm{X}}$ first and reinserting the solution resulting in [11]
where ${\bm{P}}_{\bm{H}}={\bm{H}}({\bm{H}}^{\mathrm{H}}{\bm{H}})^{-1}{\bm{H}}^{\mathrm{H}}={\bm{V}}{\bm{V}}^{\mathrm{H}}$ and $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N-J}{\bm{Y}}_{d}{\bm{Y}}_{d}^{\mathrm{H}}$, with ${\bm{Y}}_{d}$ from (2). The maximization in (56) is solved by setting ${\bm{P}}_{\bm{H}}$ equal to $\hat{{\bm{V}}}_{d}\hat{{\bm{V}}}_{d}^{\mathrm{H}}$ with $\hat{{\bm{V}}}_{d}$ holding the $J$ dominant eigenvectors of the receive sample covariance matrix $\hat{{\bm{C}}}^{(d)}_{{\bm{y}}\mid{\bm{H}}}$. This result has also been used in [26]. Additionally, it is trivial to see that the first term in (13) is minimized by ${\bm{h}}_{n}={\bm{y}}(n)$. The subspace spanned by the solution ${\bm{h}}_{n}={\bm{y}}(n)$ is the same as the subspace spanned by the $J$ eigenvectors of the sample covariance matrix $\hat{{\bm{C}}}^{(p)}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{J}{\bm{Y}}_{p}{\bm{Y}}_{p}^{\mathrm{H}}$, which ignores the additional phase information contained in the pilot observation. Thus, the overall subspace estimate $\hat{{\bm{V}}}=[{\bm{v}}_{1},\dots,{\bm{v}}_{J}]$ is found by taking the $J$ dominant eigenvectors ${\bm{v}}_{j}$ of the sample covariance matrix defined as
$\hat{{\bm{C}}}_{{\bm{y}}\mid{\bm{H}}}=\frac{1}{N}{\bm{Y}}{\bm{Y}}^{\mathrm{H}}.$ (57)
To utilize information from the previous coherence intervals, one can adaptively update the subspace using efficient tracking algorithms as proposed in, e.g., [44, 45].
IV-D Complexity Analysis
@@
The standalone GMM estimator proposed by [30] precomputes the filters used for the individual components, resulting in a complexity of $\mathcal{O}(KM^{2})$. For the standalone VAE the complexity is given as $\mathcal{O}(DM^{2})$ [32], where $D$ denotes the number of layers in the forward pass of the VAE. For our semi-blind methods, the calculation of the subspace requires $\mathcal{O}((N+J)M^{2})$. This results from calculating the sample covariance matrix with $\mathcal{O}(NM^{2})$ and taking the eigenvectors of the $J$ largest eigenvalues for the solution of (56). Using the projection approximation subspace tracking (PAST) algorithm [45], the computational complexity of calculating the subspace reduces to $\mathcal{O}(JM)$ for every update. In the case of the subspace GMM the $K$ LMMSE estimates can not be precomputed, which results in a complexity of $\mathcal{O}(K(M^{2}+JM^{2}+J^{3}))$. Similarily, the subspace VAE exhibits a complexity of $\mathcal{O}(DM^{2}+JM^{2}+J^{3})$. For the projected versions of the GMM and VAE the complexity becomes $\mathcal{O}(KM^{2}+JM^{2})$ and $\mathcal{O}(DM^{2}+JM^{2})$, respectively. One should note that the calculation for each of the $K$ components in the GMM can be parallelized. Similarly, the computations in the convolutional layers of the VAE can be parallelized, mitigating the complexity.
@@ V Baseline estimators
To compare our methods, the following baseline channel estimators are considered. Based on the found subspace $\mathrm{range}({\bm{V}})$ we can formulate the pilot-based ML estimator as $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$, which is the closed-form solution to (18). This can be interpreted as the subspace-adjusted version of the conventional least squares (LS) channel estimator given as $\hat{{\bm{h}}}_{\text{LS}}={\bm{y}}_{p}$.
Another estimator is based on the sample covariance matrix, which we can compute from the training data set ${\mathcal{H}}$ to infer the global statistics of the channels as
@@
We can use the matrix ${\bm{C}}$ as statistical prior to parameterize the semi-blind channel estimators outlined in Section III-A and Section III-B as
@@
Lastly, we compare our proposed methods to two iterative algorithms optimizing the ML formulation in (13), namely the EM from [16] and a MP variant similar to [17], which we run both until convergence or $500$ iteration, whatever comes first.

### v2

@@ IV-A GMM-based Semi-blind Channel Estimation
Algorithm 1 Subspace GMM Channel Estimator
Offline Training Phase
1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the GMM with the EM algorithm, cf. [30]
3:
[…]
8:for $j=1,\dots,J$ do
9:  ${\bm{y}}^{\prime}\leftarrow{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:  for $k=1,\dots,K$ do
12:  end for
14:end for
15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
Algorithm 2 Projected GMM Channel Estimator
Offline Training Phase
1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
2:Fit the GMM with the EM algorithm, cf. [30]
3:
[…]
8:for $j=1,\dots,J$ do
9:  $\tilde{{\bm{y}}}\leftarrow{\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p,j}$
10:  for $k=1,\dots,K$ do
12:  end for
14:end for
15:return $\hat{{\bm{h}}}_{j},\forall j=1,\dots,J$
@@
where $p(k)$, $\bm{\mu}_{k}$, and ${\bm{C}}_{k}$ are the mixing coefficients, means, and covariance matrices of the $k$-th GMM component, respectively. As we consider wireless channels, each component’s mean is set to $\bm{\mu}_{k}=\bm{0}$, cf. [43]. The fitting of the components in (44) is accomplished with the well-known EM algorithm [45] based on a set $\mathcal{H}=\{{\bm{h}}_{t}\}^{T}_{t=1}$ of $T$ channel samples as training data. Based on the formulation in (44) the conditional PDF is
@@
Thus, we have a discrete latent variable in the case of a GMM, which helps us parameterize the CME. The resulting semi-blind subspace GMM can be formulated as
@@
IV-B VAE-based Semi-blind Channel Estimation
@@
[…]
@@
[…]
@@
[…]
The resulting semi-blind VAE structure is shown in Fig. 2. In the case of a ULA or URA at the BS, the channel covariance matrix is either Toeplitz or block-Toeplitz, respectively. As shown in [43], the conditional covariance matrix at the VAE’s output preserves this structure. Thus, we parameterize the output covariance matrix as
@@
where ${\bm{Q}}={\bm{Q}}_{M}$ or ${\bm{Q}}={\bm{Q}}^{\prime}_{N_{v}}\otimes{\bm{Q}}^{\prime}_{N_{h}}$, respectively, where ${\bm{Q}}_{M}$ is a DFT matrix of size $M$ resulting in a circulant approximation, cf. [32], and ${\bm{Q}}^{\prime}_{N_{x}}$ contains the first $N_{x}$ columns of the $2N_{x}\times 2N_{x}$ DFT matrix resulting in a block-Toeplitz parameterization, cf. [47]. Further, for the (block-)Toeplitz parameterization, we can set $\bm{\mu}_{\bm{\theta}}({\bm{z}})=\bm{0}$, cf. [43].
@@
1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
@@
[…]
@@
12:     $\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}})$
@@
1:Training dataset $\mathcal{H}=\{{\bm{h}}_{t}\}_{t=1}^{T}$
@@
[…]
@@
12:     $\times(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}}))$
@@
After successfully training the VAE, the output is a local parameterization of $p({\bm{h}})$ as conditionally Gaussian
${\bm{h}}\mid{\bm{z}}\sim p_{\bm{\theta}}({\bm{h}}\mid{\bm{z}}).$ (56)
As analyzed in [32], it is a reasonable approximation to set
Based on this parameterization, we can formulate the semi-blind VAE-based estimators as
$\hat{{\bm{h}}}_{\text{proj. VAE}}=\;$ $\bm{\mu}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\bm{\theta}}({\bm{z}})\left({\bm{C}}_{\bm{\theta}}({\bm{z}})+{\bm{C}}_{\tilde{{\bm{n}}}}\right)^{-1}(\tilde{{\bm{y}}}-\bm{\mu}_{\bm{\theta}}({\bm{z}})),$ (58)
and
$\hat{{\bm{h}}}_{\text{sub. VAE}}=\;$ ${\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}\left({\bm{V}}^{\mathrm{H}}{\bm{C}}_{\bm{\theta}}({\bm{z}}){\bm{V}}+\sigma^{2}\mathbf{I}_{J}\right)^{-1}$
$\times({\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}-{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}))-{\bm{V}}{\bm{V}}^{\mathrm{H}}\bm{\mu}_{\bm{\theta}}({\bm{z}}).$ (59)
The respective estimators are summarized in Algorithm 3 and Algorithm 4. For a more detailed introduction into the VAE framework and its usage for CME parameterization, we refer the reader to [32].
(a) QPSK
(b) 16QAM
(c) Gaussian Symbols
Figure 3: NMSE over the SNR for given channel estimations in a $J=8$ user and $M=64$ antennas scenario based on $N=200$ observations, including $N_{p}=J$ pilots (if not stated otherwise) and (a) QPSK, (b) 16QAM, and (c) Gaussian data symbols, using measurement data (Sec. II-B).
IV-C Complexity Analysis
@@
The standalone GMM estimator proposed by [30] precomputes the filters used for the individual components, resulting in a complexity of $\mathcal{O}(KM^{2})$. For the standalone VAE, the complexity is given as $\mathcal{O}(DM^{2})$ [32], where $D$ denotes the number of layers in the VAE’s forward pass. For our semi-blind methods, the subspace calculation requires $\mathcal{O}((N+J)M^{2})$. This results from calculating the sample covariance matrix with $\mathcal{O}(NM^{2})$ and taking the eigenvectors of the $J$ largest eigenvalues for the solution of (34). Using the projection approximation subspace tracking (PAST) algorithm [42], the computational complexity of the subspace calculation reduces to $\mathcal{O}(JM)$ for every update. In the case of the subspace GMM, the $K$ LMMSE estimates can not be precomputed, which results in a complexity of $\mathcal{O}(K(M^{2}+JM^{2}+J^{3}))$. Similarily, the subspace VAE exhibits a complexity of $\mathcal{O}(DM^{2}+JM^{2}+J^{3})$. For the projected versions of the GMM and VAE the complexity becomes $\mathcal{O}(KM^{2}+JM^{2})$ and $\mathcal{O}(DM^{2}+JM^{2})$, respectively. One should note that the calculation for each of the $K$ components in the GMM can be parallelized. Similarly, the computations in the VAE’s convolutional layers can be parallelized, mitigating the complexity.
@@ V Baseline estimators
The following baseline channel estimators are considered for comparison with our methods. Based on the found subspace $\mathrm{range}({\bm{V}})$ we can formulate the pilot-based ML estimator as $\hat{{\bm{h}}}_{\text{ML}}={\bm{V}}{\bm{V}}^{\mathrm{H}}{\bm{y}}_{p}$, which is the closed-form solution to (16). This can be interpreted as the subspace-adjusted version of the conventional least squares (LS) channel estimator given as $\hat{{\bm{h}}}_{\text{LS}}={\bm{y}}_{p}$.
Another estimator is based on the sample covariance matrix, which we can compute from the training data set $\mathcal{H}$} to infer the channels’ global statistics as
@@
We can use the matrix ${\bm{C}}_{s}$ as a statistical prior to parameterize the semi-blind channel estimators outlined in Section III-A and Section III-B as
@@
To evaluate the gain achieved by the semi-blind adaptions of the CGLM channel estimators, we also compare to the pilot-only GMM [30] and VAE [32].
Lastly, we compare our proposed methods to two iterative algorithms optimizing the ML formulation in (13), namely the EM from [15] and a MP variant similar to [16], which we run both until convergence or $500$ iterations, whichever comes first.